OpenStudy (anonymous):

theta=3pi/4. Find in rectangular form?

OpenStudy (primeralph):

No r value?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

would the answer be tan(3pi/4)=-1

OpenStudy (primeralph):

@Directrix Please do this.

Directrix (directrix):

theta=3pi/4 inverse tangent of (y/x) = 3pi/4 y/x = tan(3pi/4) y/x = -1 Therefore, the rectangular equation for the polar equation theta=3pi/4 will be: y = -x --> a line @mutantkiller Check my work and yours, too. I used this calculation you did above: tan(3pi/4)=-1

OpenStudy (primeralph):

@Directrix I thought of that too as an infinite line, but for some reason, I expected a ray.

OpenStudy (anonymous):

i have one other question

OpenStudy (anonymous):

how do you change x+y=2 into polar form?

Directrix (directrix):

@primeralph I pondered that same question. I have seen theta = some angle graphed as a ray and then as that same ray with the ray opposite it, forming a line. l I don't know that there is clarity on that issue. I need to do some reading.

OpenStudy (primeralph):

@Directrix Sometimes the basic stuff gets us too. @mutantkiller Just put x =rcostheta and y =rsinetheta and simplify if possible.

OpenStudy (anonymous):

ok thanks

OpenStudy (primeralph):

You're welcome.